you wish to replenish your stock of dishes by purchasing 250 sets for the restaurant. You have two different dish designs available. one cost $20 per set and the other is $45 per set. If you have $6800 to spend , how many of each design should you order? x=$20 dishes and y=$45 dishes.
e-mail natalieloftin46@yahoo.com

Answer by rfer(16322)

x+y=250
20x+45y=6800
20(250-y)+45y=6800
5000-20y+45y=6800
25y=1800
y=72
-------------
x+72=250
x=178